# Integrating Spatiotemporal Features in LSTM for Spatially Informed COVID-19 Hospitalization Forecasting

### v1

@@ 5 Discussion
To improve our model’s performance, we also tested an additional feature based on Google Search Trends as an early signal of hospitalizations. We experimented with two groups of search trends, one including COVID-related terms (e.g., “COVID-19”, “Coronavirus”, and “COVID Vaccine”) and one narrowed down to COVID-related symptoms (e.g., “loss of taste”, “loss of smell”, “cough”, and “fever”). Though both COVID-related terms and symptoms show high correlations with the cases and hospitalizations, we found that incorporating Google search trends in our models only improves performance marginally during the increase period and deteriorates performance over the peak and decrease periods. We also found that COVID-related symptom search is more helpful than COVID-related terms during the increasing phase. One of the potential reasons that COVID-related symptoms did not improve over the peak and decreasing period may be the overlap with the flu season. COVID-19 and flu share similar symptoms like fever, cough, runny nose, and other symptoms, which might not correlate highly with COVID-19 cases and hospitalizations, such as ageusia, anosmia, and pneumonia. Given that the gain in performance was marginal and only for a subset of our evaluation weeks, we decided not to incorporate this feature in the final models and results presented. Furthermore, the lack of an official API for accessing Google Search Trends, coupled with the inefficiency of using third-party tools like pytrends for real-time forecasting due to their slow performance, reinforced our decision not to include this feature in the final models and subsequent results.
The improvements made by our model in predicting COVID-19 hospitalizations using SCI-derived features and a novel stacked architecture of LSTMs underscores its potential applicability to other infectious diseases with similar transmission dynamics and data collection practices. The inherent flexibility of our deep learning framework, which is capable of integrating diverse types of spatiotemporal data (e.g., county/state-level; daily/weekly), makes it well-suited for diseases like influenza that exhibit seasonal patterns and are influenced by human mobility and social behavior. The FluSight collaborative forecasting Hub, organized by the US CDC, was modeled after the COVID-19 Forecast Hub, and offers an ideal benchmark for testing the model’s generalizability to influenza (Centers for Disease Control and Prevention 2024, Reich et al. 2019). Future research will aim to evaluate the model’s efficacy with data from other diseases.
@@ 6 Conclusion
In this paper, we presented a novel parallel Long Short-Term Memory (LSTM) framework with spatiotemporal features derived from Facebook’s Social Connectedness Index to capture spatial spillover effect. Our parallel architecture learns and combined short- and long-term temporal dependencies effectively. We also presented a novel multi-horizon ensembling strategy to balance between forecast consistency, uncertainty, and performance. We evaluated our forecasts on hospitalization reports during the Omicron surge and found that our approach outperforms multiple baselines over the 28-day horizon on average, with high improvement margin towards the end. Long-term forecast, and the biggest average improvement on $28^{th}$ day forecast can reach about 70 incident hospitalizations per state. Our evaluation results also demonstrated that our spatiotemporal feature, SPH, improved forecasts and reduced model variance. It is worth reiterating that the Omicron wave was the most challenging period for hospitalization forecasting models. In previous waves of the virus, hospitalizations followed case surges with a rough two to three-week delay. With the change of coupling dynamics between cases and hospitalization cases by the Omicron variant that led to an explosive number of cases but fewer hospitalizations, the improvements achieved by our approach over baselines are notable. The strong predictive power of our approach is attributable to the novel spatiotemporal features integrated in LSTM, use of incidence rates to decouple from population size, the parallel LSTM architecture with learnable balancing of short- and long-term dependencies, a novel data-driven, a multi-horizon ensembling strategy.
@@
The COVID-19 pandemic and the underperformance of forecasting models on the Forecast Hub (Lopez et al. 2024) shows the need for future research in this area. Despite the fact that statistical approaches have shown high forecasting skills, less than one-third of teams contributing to the Hub used statistical models. One potential reason is the legacy of compartmental models and their key parameters, like $R_{0}$. Explainability and interoperability are valuable for epidemiologists to ensure the reliability of forecasting and to interpret the dynamics of epidemics. Therein lies the gap between the forecasting skill of statistical methods and the explainability of compartmental methods, and future research should investigate ways of better combining the strengths of those families of models.
@@
Another worthwhile direction lies in the use of spatiotemporal graph neural networks (GNNs). In our approach, the spatiotemporal feature that captures spatiotemporal spillover was formulated. In a GNN, nodes can represent spatial regions, while inter-region connectedness can be used as edge features (Kapoor et al. 2020). GNNs have proved capable of node and graph classification tasks (Hamilton et al. 2017). However, their performance for regression tasks such as those required by the Forecasting Hub can be negatively affected due to oversmoothing in GNNs (Rusch et al. 2023), where further research is warranted.
@@ 7 Data and Codes Availability Statement
Zhongying Wang: Data curation; Software; Methodology; Writing – original draft; Visualization; Formal analysis.
Thoai D. Ngo: Conceptualization; Funding acquisition; Writing – review & editing; Validation.
Benjamin Lucas: Methodology consultation; Writing – review & editing.
Hamidreza Zoraghein: Data curation; Writing – review & editing.
Morteza Karimzadeh: Conceptualization; Methodology; Supervision; Funding acquisition; Project administration; Writing – original draft; Writing – review & editing.

### v2

@@ 5 Discussion
To improve our model’s performance, we also tested an additional feature based on Google Search Trends as an early signal of hospitalizations. We experimented with two groups of search trends, one including COVID-related terms (e.g., “COVID-19”, “Coronavirus”, and “COVID Vaccine”) and one narrowed down to COVID-related symptoms (e.g., “loss of taste”, “loss of smell”, “cough”, and “fever”). Though both COVID-related terms and symptoms showed high correlations with the cases and hospitalizations, we found that incorporating Google search trends in our models only improved performance marginally during the increase period and deteriorated performance over the peak and decrease periods. We also found that COVID-related symptom search was more helpful than COVID-related terms during the increasing phase. One of the potential reasons that COVID-related symptoms did not improve over the peak and decreasing period may be the overlap with the flu season. COVID-19 and flu share similar symptoms, such as fever, cough, runny nose, and other symptoms, which may not correlate highly with COVID-19 cases and hospitalizations, such as ageusia, anosmia, and pneumonia. Given that the performance gain was marginal and only observed for a subset of our evaluation weeks, we decided not to incorporate this feature into the final models and results presented. Furthermore, the lack of an official API for accessing Google Search Trends, coupled with the inefficiency of using third-party tools such as pytrends for real-time forecasting due to their slow performance, reinforced our decision not to include this feature in the final models and subsequent results.
The improvements made by our model in predicting COVID-19 hospitalizations using SCI-derived features and a novel stacked architecture of LSTMs underscore its potential applicability to other infectious diseases with similar transmission dynamics and data collection practices. The inherent flexibility of our deep learning framework, which is capable of integrating diverse types of spatiotemporal data (e.g., county/state-level; daily/weekly), makes it well-suited for diseases like influenza that exhibit seasonal patterns and are influenced by human mobility and social behavior. The FluSight collaborative forecasting Hub, organized by the US CDC, was modeled after the COVID-19 Forecast Hub, and offers an ideal benchmark for testing the model’s generalizability to influenza (Centers for Disease Control and Prevention 2024, Reich et al. 2019). Future research will aim to evaluate the model’s efficacy with data from other diseases.
@@ 6 Conclusion
In this paper, we presented a novel parallel Long Short-Term Memory (LSTM) framework with spatiotemporal features derived from Facebook’s Social Connectedness Index to capture the spatial spillover effect. Our parallel architecture learns and combines short- and long-term temporal dependencies effectively. We also presented a novel multi-horizon ensembling strategy to balance between forecast consistency, uncertainty, and performance. We evaluated our forecasts on hospitalization reports during the Omicron surge and found that our approach outperforms multiple baselines over the 28-day horizon on average, with a high improvement margin towards the end. Long-term forecast, and the biggest average improvement on the $28^{th}$ day forecast can reach about 70 incident hospitalizations per state. Our evaluation results also demonstrated that our spatiotemporal feature, SPH, improved forecasts and reduced model variance. It is worth reiterating that the Omicron wave was the most challenging period for hospitalization forecasting models. In previous waves of the virus, hospitalizations typically followed case surges with a delay of approximately two to three weeks. With the change in coupling dynamics between cases and hospitalizations caused by the Omicron variant, which led to an explosive number of cases but fewer hospitalizations, the improvements achieved by our approach over baselines were notable. The strong predictive power of our approach is attributable to the novel spatiotemporal features integrated in LSTM, the use of incidence rates to decouple from population size, the parallel LSTM architecture with learnable balancing of short- and long-term dependencies, and a novel data-driven, multi-horizon ensembling strategy.
@@
The COVID-19 pandemic and the underperformance of forecasting models on the Forecast Hub (Lopez et al. 2024) show the need for future research in this area. Despite the fact that statistical approaches have demonstrated high forecasting skills, fewer than one-third of the teams contributing to the Hub utilized statistical models. One potential reason is the legacy of compartmental models and their key parameters, such as $R_{0}$. Explainability and interoperability are valuable for epidemiologists to ensure the reliability of forecasting and to interpret the dynamics of epidemics. Therein lies the gap between the forecasting skill of statistical methods and the explainability of compartmental methods; future research should investigate ways to better combine the strengths of these two families of models.
@@
Another worthwhile direction lies in the use of spatiotemporal graph neural networks (GNNs). In our approach, the spatiotemporal feature that captures spatiotemporal spillover was formulated. In a GNN, nodes can represent spatial regions, while inter-region connectedness can be used as edge features (Kapoor et al. 2020). GNNs have demonstrated their capability in node and graph classification tasks (Hamilton et al. 2017). However, their performance on regression tasks, such as those required by the Forecasting Hub, can be negatively affected due to oversmoothing in GNNs (Rusch et al. 2023), indicating the need for further research.
@@ 7 Data and Codes Availability Statement
Zhongying Wang is a graduate researcher in the Department of Geography at the University of Colorado Boulder, USA (E-mail: zhongying.wang@colorado.edu). His research focuses on geospatial data science and machine learning for public health. Wang led the data curation, software implementation, methodology design, formal analysis, visualization, and drafting of the manuscript.
Thoai D. Ngo is Professor in the Heilbrunn Department of Population and Family Health, Columbia University Mailman School of Public Health, USA (E-mail: tn2571@cumc.columbia.edu). His work focuses on global public health and epidemiological research. Ngo conceived the study, secured funding, validated the analysis, and critically revised the manuscript.
Benjamin Lucas is a Post-doctoral Associate in the Department of Geography at the University of Colorado Boulder, USA (E-mail: ben.lucas@nau.edu). His interests include machine learning for remote-sensing applications. Lucas advised on the study methodology and contributed to the review and editing of the manuscript.
Hamidreza Zoragheinis a Research Scientist at the Population Council, New York, USA (E-mail: hrz1365@gmail.com). He specialises in spatial data analysis and public-health applications. Zoraghein curated data resources and participated in the critical revision of the manuscript.
Morteza Karimzadeh: is Assistant Professor of Geography at the University of Colorado Boulder, USA (E-mail: karimzadeh@colorado.edu). His research integrates geospatial data science, remote sensing and deep learning for public-health, social and environmental applications. Karimzadeh originated the study concept, directed the methodology, supervised the research team, oversaw project administration and funding acquisition, and contributed to both the original drafting and critical revision of the manuscript.